# Identifying Astrophysical Anomalies in 99.6 Million Source Cutouts from the Hubble Legacy Archive Using AnomalyMatch

### v1

@@ 5 Discussion
Other works focused on anomaly detection often must begin from using similarity searches, to citizen scientists to identify large samples for further searches (Lintott et al., 2008; Lochner & Bassett, 2021; Mantha et al., 2024; Euclid Collaboration et al., 2025b). Conversely, AnomalyMatch can be used as a baseline approach to grow existing samples of anomalies from the literature. It can also be employed in a flexible manner, where other anomalies not included in the initial labelled dataset can be detected and extracted during the hunt for other objects.
@@
In this work, we find 811 unreferenced objects in our sample of 1,339 objects. However, by the nature of using the HLA and, specifically, HST to search for anomalies, we are limiting our capability for identifying new objects. The HST is not a survey telescope: observations must be applied for, and PIs specifically request the coordinates and objects they wish to observe. Therefore, any anomalies we do find are either in the background of observations of another astrophysical source of interest, or are the target of the observation in question. The latter is particularly true for edge-on proto-planetary disks, where these targets of exceptional interest are rarely observed ‘by accident’ in the background of targeted observations. The fact that $\approx 65\%$ of our anomalies are not represented in the literature thus shows the exceptional potential of AnomalyMatch for application to survey telescopes’ data.
@@
In the future, AnomalyMatch will be used to search for specific types of anomalies that the user can hone in on. By ranking objects in this way, we will be able to rapidly find and increase the size of our samples of different anomalies while using visual inspection. As found in Gomez & O’Ryan (2025), anomalies identified with the 1% highest scores have a high precision, i.e. chance to be only or almost only anomalies - assuming such objects are present in the data of them in the first place. However, extracting them still requires visual inspection by the user.
@@ 6 Conclusion
We have developed and deployed an innovative approach based on semi-supervised and active learning for the purpose of identifying sources with anomalous morphologies. The method - AnomalyMatch - has been created as an out-of-the-box tool, which has been integrated into the ESA Datalabs platform. It has an intuitive and interactive UI, which seamlessly combines semi-supervised learning methods with active learning so a user can identify objects of interest efficiently, and with few labels. The full benchmarking and testing of this algorithm is described in Gomez & O’Ryan (2025).
@@
Upon completion of development, and training this model and applying it across the HLA, we selected the top 1,339 de-duplicated sources. We then classified these into different sub-classes of anomalies based on morphology. These included 629 mergers, 219 gravitational lenses, 35 jellyfish galaxies and 2 edge-on protoplanetary disks. Using SIMBAD and ESASky, we conducted a literature search based on the coordinates of each object, and to investigate if they had been included in other catalogues and samples. We find that $\approx 65\%$, 811, of the objects do not appear in the literature.
@@
Acknowledgements
The authors would like to acknowledge and thank Chris Evans, Sandor Kruk, Maria Teresa Nardone, Pedro Mas Buitrago, and Laslo Ruhberg for their valuable feedback on this manuscript. DOR acknowledges the support of the ESA Research Fellowship in Space Science program. This work was published as part of his research fellowship.
This work makes extensive use of datasets from the Hubble Legacy Archive. Stored here are observations with the NASA/ESA Hubble Space Telescope, which is a collaboration between the Space Telescope Science Institute (STScI / NASA), The Space Telescope European Coordinating Facility (ST-ECF/ESA) and the Canadian Astronomy Data Centre (CADC/NRC/CSA). This work also made extensive use of theHubble Source Cataloue, derived from the HLA. This is fully described in Whitmore et al. (2016). To efficiently interface with the HLA, ESA’s science platform ESA Datalabs was used. ESA Datalabs (data.esa.int) is an initiative by ESA’s Data Science and Archives division in the Science and Operations Department, Directorate of Science. Navarro et al. (2024) describes the platform.
To study the objects we found, and aid in their classifications, extensive use of the ESASky platform and SIMBAD database were used. The ESASky platform was developed by the ESAC Science Data Centre (ESDC) team and maintained alongside other ESA science mission’s archives at ESA’s European Space Astronomy Centre (ESAC, Madrid, Spain). A full description of the platform can be found in Baines et al. (2017) and Giordano et al. (2018). This research has also made use of the SIMBAD database, operated at CDS, Strasbourg France. It is fully described in Wenger et al. (2000).
This research made use of many open-source Python packages and scientific computing systems. These included Matplotlib Hunter (2007), Pandas (McKinney, 2010),and numpy (Harris et al., 2020). This work also extensively used the community-driven Python package Astropy (Astropy Collaboration et al., 2018, 2022).
This work made extensive use of the new tool AnomalyMatch, described in the companion paper Gomez & O’Ryan (2025). For writing and editing, no additional AI tools were used in this manuscript.

### v3

@@ 5 Discussion
Other works focused on anomaly detection often must begin from using similarity searches, to citizen scientists to identify large samples for further searches (Lintott et al., 2008; Lochner & Bassett, 2021; Mantha et al., 2024; Euclid Collaboration et al., 2025c). Conversely, AnomalyMatch can be used as a baseline approach to grow existing samples of anomalies from the literature. It can also be employed in a flexible manner, where other anomalies not included in the initial labelled dataset can be detected and extracted during the hunt for other objects.
@@
In this work, we find 810 unreferenced objects in our sample of 1,338 objects. However, by the nature of using the HLA and, specifically, HST to search for anomalies, we are limiting our capability for identifying new objects. The HST is not a survey telescope: observations must be applied for, and PIs specifically request the coordinates and objects they wish to observe. Therefore, any anomalies we do find are either in the background of observations of another astrophysical source of interest, or are the target of the observation in question. The latter is particularly true for edge-on proto-planetary disks, where these targets of exceptional interest are rarely observed ‘by accident’ in the background of targeted observations. The fact that $\approx 65\%$ of our anomalies are not represented in the literature thus shows the exceptional potential of AnomalyMatch for application to survey telescopes’ data.
@@
In the future, AnomalyMatch will be used to search for specific types of anomalies that the user can hone in on. By ranking objects in this way, we will be able to rapidly find and increase the size of our samples of different anomalies while using visual inspection. As found in Gómez & O’Ryan (2025), anomalies identified with the 1% highest scores have a high precision, i.e. chance to be only or almost only anomalies - assuming such objects are present in the data of them in the first place. However, extracting them still requires visual inspection by the user.
@@ 6 Conclusion
We have developed and deployed an innovative approach based on semi-supervised and active learning for the purpose of identifying sources with anomalous morphologies. The method - AnomalyMatch - has been created as an out-of-the-box tool, which has been integrated into the ESA Datalabs platform. It has an intuitive and interactive UI, which seamlessly combines semi-supervised learning methods with active learning so a user can identify objects of interest efficiently, and with few labels. The full benchmarking and testing of this algorithm is described in Gómez & O’Ryan (2025).
@@
Upon completion of development, and training this model and applying it across the HLA, we selected the top 1,338 de-duplicated sources. We then classified these into different sub-classes of anomalies based on morphology. These included 629 mergers, 140 gravitational lenses, 37 jellyfish galaxies and 2 edge-on protoplanetary disks. Using SIMBAD and ESASky, we conducted a literature search based on the coordinates of each object, and to investigate if they had been included in other catalogues and samples. We find that $\approx 65\%$, 811, of the objects do not appear in the literature.
@@
Acknowledgements.
The authors would like to acknowledge and thank Chris Evans, Sandor Kruk, Maria Teresa Nardone, Pedro Mas Buitrago, and Laslo Ruhberg for their valuable feedback on this manuscript. DOR acknowledges the support of the ESA Research Fellowship in Space Science program. This work was published as part of his research fellowship. The authors also acknowledge and thank the thorough review of this work by the external referee. Their reports strengthened this work significantly. This work makes extensive use of datasets from the Hubble Legacy Archive. Stored here are observations with the NASA/ESA Hubble Space Telescope, which is a collaboration between the Space Telescope Science Institute (STScI / NASA), The Space Telescope European Coordinating Facility (ST-ECF/ESA) and the Canadian Astronomy Data Centre (CADC/NRC/CSA). This work also made extensive use of theHubble Source Cataloue, derived from the HLA. This is fully described in Whitmore et al. (2016). To efficiently interface with the HLA, ESA’s science platform ESA Datalabs was used. ESA Datalabs (data.esa.int) is an initiative by ESA’s Data Science and Archives division in the Science and Operations Department, Directorate of Science. Navarro et al. (2024) describes the platform. To study the objects we found, and aid in their classifications, extensive use of the ESASky platform and SIMBAD database were used. The ESASky platform was developed by the ESAC Science Data Centre (ESDC) team and maintained alongside other ESA science mission’s archives at ESA’s European Space Astronomy Centre (ESAC, Madrid, Spain). A full description of the platform can be found in Baines et al. (2017) and Giordano et al. (2018). This research has also made use of the SIMBAD database, operated at CDS, Strasbourg France. It is fully described in Wenger et al. (2000). This research made use of many open-source Python packages and scientific computing systems. These included Matplotlib Hunter (2007), Pandas (McKinney, 2010),and numpy (Harris et al., 2020). This work also extensively used the community-driven Python package Astropy (Astropy Collaboration et al., 2018, 2022). This work made extensive use of the new tool AnomalyMatch, described in the companion paper Gómez & O’Ryan (2025). For writing and editing, no additional AI tools were used in this manuscript.